OpenStudy (anonymous):

Given V = 2πr2h, solve for h

OpenStudy (anonymous):

V = 2(pi)r^2h (divide by 2(pi)r^2) V/2(pi)r^2 = h

mathslover (mathslover):

\[\large{ V = 2\pi r^2 h}\] \[\large{\frac{V}{2\pi r^2} = h}\]

OpenStudy (aravindg):

Hi, I can see that you are a new member here at OpenStudy so, I would first like to say Welcome to OpenStudy, I would like to point you to the chat pods these chat pods are where you can make new friends and talk to new people just like you, I would also like to emphasize our "NON-CHEATING POLICY" we ask here at OpenStudy that you do NOT post exam/Test questions if you are caught doing this we will notify your current school, I would also like to ask you to check out the OpenStudy "Code of Conduct" http://openstudy.com/code-of-conduct Please read this carefully and thoroughly. Welcome to the OpenStudy Community! Ambassador